Please can you help me with this word problem, it will be appreciated. Brandon is putting together a rectangular fence around his garden to keep the animals from stealing his vegetables. The perimeter of the rectangle is 72 feet. The ratio of the width to the length is 5 to 7. What are the dimensions of the rectangle?
Found 2 solutions by ikleyn, ankor@dixie-net.com:
Answer by ikleyn(52803)

.
15 and 21.


Answer by ankor@dixie-net.com(22740)

The perimeter of the rectangle is 72 feet.
The ratio of the width to the length is 5 to 7.
What are the dimensions of the rectangle?
:
let x = the multiplier
then
5x = the width
and
7x = the length
:
The perimeter = 72
2(7x) + 2(5x) = 72
simplify, divide by 2
7x + 5x = 36
12x = 36
x = 36/12
x =3 is the multiplier
therefore
7(3) = 21 ft is the length
and
5(3) = 15 ft is the width
:
:
Check
2(21) + 2(15) = 72